Brook’s Window Shields Inc. is trying to calculate its cost of capital for use in a capital budgeting decision. Mr. Glass, the vice-president of finance, has given you the following information and has asked you to compute the weighted average cost of capital. The company currently has outstanding a bond with a 11.2 percent coupon rate and another bond with a 7.2 percent coupon rate. The firm has been informed by its investment banker that bonds of equal risk and credit rating are now selling to yield 12.2 percent. The common stock has a price of $70 and an expected dividend (D1) of $5.50 per share. The firm's historical growth rate of earnings and dividends per share has been 7.2 percent, but security analysts on Wall Street expect this growth to slow to 5 percent in future years. The preferred stock is selling at $66 per share and carries a dividend of $6.50 per share. The corporate tax rate is 35 percent. The flotation cost is 3.1 percent of the selling price for preferred stock. The optimal capital structure is 30 percent debt, 25 percent preferred stock, and 45 percent common equity in the form of retained earnings. Compute the cost of capital for the individual components in the capital structure. then Calculate the weighted cost of each source of capital and the weighted average cost of capital.
Answer #1

kD = 12.2%

kE = [D1/P0] + g = [$5.50/$70] + 0.072 = 0.0786 + 0.072 = 0.1506, or 15.06%

kP = Annual Dividend / [Current Price x (1 - fc)]

= $6.50 / [$66 x (1 - 0.031)] = 0.1016, or 10.16%

WACC = [wD x kD x (1 - t)] + [wP x kP] + [wE x kE]

= [0.30 x 12.2% x (1 - 0.35)] + [0.25 x 10.16%] + [0.45 x 15.06%] = 2.38% + 2.54% + 6.78% = 11.70%